A woodcutter determines the height of a tall tree by first measuring a smaller one 125 ft away, then moving so that his eyes are in the line of sight along the tops of the trees, and measuring how far he is standing from the small tree. Suppose the small tree is 20 ft tall, the man is 25 ft from the small tree, and his eye level is 5 ft above the ground. How tall is the taller tree?
Answer by scott8148(6628)

the height (ABOVE EYE LEVEL) is proportional to the distance to the trees (similar triangles)

25 / (20-5) = (125+25) / (x-5)

"cross" multiplying ___ 25x - 125 = 2250 ___ 25x = 2375 ___ x = 95